# Characterizing Kadison–Schwarz maps on $M_{3}$

### v2

@@ IV.2 Proof of Theorem 1
For completely positive maps the Kadison–Schwarz inequality is automatically satisfied. Theorem 1 shows that, on $M_{3}$, the Kadison–Schwarz property may hold under substantially weaker assumptions, which can be expressed explicitly in terms of the Bloch parameters.
V Conclusions
In this work we analyzed the Kadison–Schwarz (KS) property for unital linear maps acting on matrix algebras, using the Bloch–Gell–Mann representation. After reviewing general properties of KS maps on $M_{d}$, we specialized the analysis to the case of $M_{3}$, where the explicit structure of the Lie algebra $\mathfrak{su}(3)$ allows for analytic control of operator products appearing in the Kadison–Schwarz inequality.
Our main result provides an explicit sufficient condition for the Kadison–Schwarz property of unital maps on $M_{3}$, formulated directly in terms of the parameters of the Bloch matrix. In contrast to approaches based on numerical optimization or on complete positivity criteria, the condition derived here is purely analytic and exploits structural features of the $\mathfrak{su}(3)$ algebra. In particular, the cancellation of all contributions involving the antisymmetric structure constants and the resulting reduction to estimates involving only the symmetric constants play a central role.
The results clarify the position of KS maps between positive and completely positive maps on $M_{3}$. They demonstrate that the Kadison–Schwarz property may hold under assumptions substantially weaker than complete positivity, while still imposing nontrivial constraints beyond mere positivity. This highlights the usefulness of KS maps as an intermediate class, relevant in contexts where complete positivity is not required but additional structure is desirable.
The framework developed in this work naturally extends to arbitrary dimension $d$. The Bloch representation and the reduction via unitary equivalence remain valid on $M_{d}$, and the Kadison–Schwarz property is invariant under changes of the generator basis. However, for $d>3$ the control of expressions involving the symmetric structure constants of $\mathfrak{su}(d)$ becomes increasingly subtle, and the derivation of sharp analytic bounds requires additional tools. A systematic extension of the present analysis to higher dimensions, including dimension-dependent estimates for KS maps on $M_{d}$, is left for future work.
We expect that the methods presented here may also be useful in the study of Kadison–Schwarz properties of dynamical maps and generators appearing in the theory of open quantum systems, as well as in the analysis of intermediate classes of positive maps beyond complete positivity.
@@ A.1 Structure constants of $\mathfrak{su}(3)$
where $f_{ijk}$ and $d_{ijk}$ are the antisymmetric and symmetric structure constants of $\mathfrak{su}(3)$, respectively. Explicit values of these constants can be found, e.g., in Refs. [Georgi1999, Slansky].
@@ A.2 Norm estimates
where the coefficients $y_{k}$ arise as quadratic combinations of the Bloch components $w_{i}$ and the parameters $\mu_{j}$. Using the submultiplicativity of the operator norm and the orthonormality of the Gell–Mann basis, one obtains the bound
$\|Y\|\;\leq\;\sum_{k=1}^{8}|y_{k}|\,\|\lambda_{k}\|=\sqrt{2}\,\sum_{k=1}^{8}|y_{k}|.$ (18)
The coefficients $y_{k}$ contain terms of the form
$\sum_{i,j}(\mu_{i}-\mu_{j})\,d_{ijk}\,w_{i}\overline{w_{j}}.$
Applying the Cauchy–Schwarz inequality and using the finiteness of the set of nonvanishing symmetric structure constants, one finds
$\sum_{k=1}^{8}|y_{k}|\;\leq\;C_{3}\,\max_{i,j}|\mu_{i}-\mu_{j}|\,\|X^{\dagger}X\|,$ (19)
where the constant $C_{3}>0$ depends only on the numerical values of the coefficients $d_{ijk}$.
Combining (18) and (19) yields the estimate
$\biggl\|\sum_{k=1}^{8}y_{k}\lambda_{k}\biggr\|\;\leq\;C_{3}\,\max_{i,j}|\mu_{i}-\mu_{j}|\,\|X^{\dagger}X\|,$ (20)
which is precisely the bound stated in Eq. (15).
@@ A.3 Remarks
The constant $C_{3}$ is not universal and depends on the chosen normalization of the generators. Its explicit numerical value is irrelevant for the formulation of Theorem 1. What matters is that $C_{3}$ is finite and depends only on the algebraic structure of $\mathfrak{su}(3)$, rather than on the specific map $\Phi$. This ensures that the condition (13) is purely algebraic and independent of the particular input operator $X$.

### v3

@@ V Illustrative diagonal family
Theorem 1 is formulated for unital positive maps. Accordingly, in the present example the Kadison–Schwarz condition is interpreted only within the parameter region where the map $\Phi_{t}$ is positive, namely $\tfrac{1}{2}\leq t\leq 1$. Although the spectral condition of the theorem is trivially satisfied for all $t$ in the isotropic case ($t=s$), values t<1/2t<-1/2 are excluded since positivity is a necessary prerequisite for the Kadison–Schwarz property.
For the same isotropic subfamily one has $\max_{i,j}|\mu_{i}-\mu_{j}|=0$, and hence Theorem 1 guarantees the Kadison–Schwarz property throughout the region where the map is positive, i.e. for $\tfrac{1}{2}\leq t\leq 1$. Consequently, within the class of positive unital maps, the sufficient KS region strictly extends beyond the completely positive regime.
The family $T(t,s)$ therefore provides a geometric illustration of how analytic bounds derived from the $\mathfrak{su}(3)$ structure control deviations from complete positivity while still ensuring the Kadison–Schwarz property.
VI Conclusions
We have derived an explicit sufficient criterion for the Kadison–Schwarz (KS) property of unital positive maps on $M_{3}$ within the Bloch–Gell–Mann representation. For diagonal Bloch matrices $T=\operatorname{diag}(\mu_{1},\dots,\mu_{8})$ satisfying the contraction bound $\max_{k}|\mu_{k}|\leq 1$, the KS condition follows whenever the spectral spread $\max_{i,j}|\mu_{i}-\mu_{j}|$ is small enough compared to the ratio $c_{3}(\mu)/C_{3}$, where $c_{3}(\mu)$ quantifies the uniform dominance of the scalar term in the KS expansion and $C_{3}$ is a structural constant determined solely by the symmetric tensor $d_{ijk}$ of $\mathfrak{su}(3)$.
A key structural feature behind the proof is that for diagonal $T$ all contributions involving the antisymmetric structure constants $f_{ijk}$ cancel in the KS expression, reducing the problem to norm estimates governed entirely by $d_{ijk}$. This yields a transparent dominance argument: the traceless fluctuation term is controlled by $\max_{i,j}|\mu_{i}-\mu_{j}|$ and is dominated by the positive scalar part whenever (15) holds. In particular, within the positive unital regime the sufficient KS region may extend strictly beyond complete positivity, illustrating that the KS property captures a genuinely intermediate level of positivity.
The present analysis suggests several directions for further work. First, it would be desirable to obtain sharper (and possibly more explicit) lower bounds on $c_{3}(\mu)$ for natural subclasses of diagonal maps, leading to more quantitative KS regions in parameter space. Second, extending the approach to $M_{d}$ with $d>3$ requires dimension-dependent control of expressions involving the symmetric structure constants of $\mathfrak{su}(d)$ and a careful treatment of unitary equivalence, since not every orthogonal transformation in $SO(d^{2}-1)$ arises from the adjoint action of $SU(d)$ (see [1, 7]). We expect that the analytic mechanism used here may be useful in the study of KS properties of dynamical maps and generators in open quantum systems, as well as in the systematic exploration of intermediate positivity classes beyond complete positivity. It would be interesting to understand whether analogous dominance mechanisms persist for broader classes of positive maps beyond the diagonal Bloch setting.
@@ A.1 Structure constants of $\mathfrak{su}(3)$
where $f_{ijk}$ and $d_{ijk}$ are the antisymmetric and symmetric structure constants of $\mathfrak{su}(3)$, respectively. Explicit values of these constants can be found, e.g., in Refs. [7, 18].
@@ A.2 Norm estimates
where the coefficients $y_{k}$ arise as quadratic combinations of the Bloch components $w_{i}$ and the parameters $\mu_{j}$. In particular, when $Y$ is the traceless part of the KS expansion (16), we have $y_{k}=\beta_{k}(X)$ for $k=1,\dots,8$. Using the submultiplicativity of the operator norm and the orthonormality of the Gell–Mann basis, one obtains the bound
$\|Y\|\;\leq\;\sum_{k=1}^{8}|y_{k}|\,\|\lambda_{k}\|\leq\Bigl(\max_{1\leq k\leq 8}\|\lambda_{k}\|\Bigr)\sum_{k=1}^{8}|y_{k}|.$ (28)
Since $\max_{k}\|\lambda_{k}\|$ is a fixed numerical constant for the chosen normalization, we incorporate it into the definition of $C_{3}$ below. The coefficients $y_{k}$ contain terms of the form
$\sum_{i,j=1}^{8}(\mu_{i}-\mu_{j})\,d_{ijk}\,w_{i}\overline{w_{j}}.$ (29)
To make the dependence explicit, we introduce the auxiliary constant
$C_{3}^{(0)}:=\sup_{\|u\|_{2}=\|v\|_{2}=1}\sum_{k=1}^{8}\left|\sum_{i,j=1}^{8}d_{ijk}\,u_{i}\overline{v_{j}}\right|,$ (30)
where $\|\cdot\|_{2}$ denotes the Euclidean norm on $\mathbb{C}^{8}$. Here $d_{ijk}\in\mathbb{R}$ for the Gell–Mann normalization. Since the set of nonvanishing symmetric structure constants $d_{ijk}$ is finite, the quantity (30) is finite and depends only on the algebraic structure of $\mathfrak{su}(3)$ and the chosen normalization of the generators.
Applying the Cauchy–Schwarz inequality to (29) and using (30), one obtains
$\sum_{k=1}^{8}|y_{k}|\;\leq\;C_{3}^{(0)}\,\max_{i,j}|\mu_{i}-\mu_{j}|\,\|w\|_{2}^{2},$ (31)
where $C_{3}^{(0)}$ depends only on the symmetric structure constants $d_{ijk}$ and the chosen normalization of the generators.
To relate $\|w\|_{2}$ to an operator norm quantity, note that the Bloch–Gell–Mann expansion $X=w_{0}\lambda_{0}+\sum_{k=1}^{8}w_{k}\lambda_{k}$ satisfies, by $\operatorname{Tr}(\lambda_{a}^{\dagger}\lambda_{b})=2\delta_{ab}$,
$\|X\|_{\mathrm{HS}}^{2}=2\bigl(|w_{0}|^{2}+\|w\|_{2}^{2}\bigr),\qquad\text{hence}\qquad\|w\|_{2}^{2}\leq\frac{1}{2}\,\|X\|_{\mathrm{HS}}^{2}.$ (32)
Moreover, for $X\in M_{3}$ one has $\|X\|_{\mathrm{HS}}\leq\sqrt{3}\,\|X\|$, and therefore
$\|w\|_{2}^{2}\leq\frac{1}{2}\,\|X\|_{\mathrm{HS}}^{2}\leq\frac{3}{2}\,\|X\|^{2}=\frac{3}{2}\,\|X^{\dagger}X\|.$ (33)
Combining (31) and (33) yields
$\sum_{k=1}^{8}|y_{k}|\;\leq\;\frac{3}{2}\,C_{3}^{(0)}\,\max_{i,j}|\mu_{i}-\mu_{j}|\,\|X^{\dagger}X\|.$ (34)
Together with (28) this yields
$\biggl\|\sum_{k=1}^{8}y_{k}\lambda_{k}\biggr\|\;\leq\;\Bigl(\max_{1\leq k\leq 8}\|\lambda_{k}\|\Bigr)\,\frac{3}{2}\,C_{3}^{(0)}\,\max_{i,j}|\mu_{i}-\mu_{j}|\,\|X^{\dagger}X\|.$ (35)
For notational simplicity we redefine
$C_{3}:=\Bigl(\max_{1\leq k\leq 8}\|\lambda_{k}\|\Bigr)\,\frac{3}{2}\,C_{3}^{(0)},$ (36)
which is still a structural constant depending only on $\mathfrak{su}(3)$ (and the chosen normalization), and we arrive at the bound stated in the main text:
$\biggl\|\sum_{k=1}^{8}y_{k}\lambda_{k}\biggr\|\;\leq\;C_{3}\,\max_{i,j}|\mu_{i}-\mu_{j}|\,\|X^{\dagger}X\|.$ (37)
@@ A.3 Remarks
The auxiliary constant $C_{3}^{(0)}$ is the operator norm of the bilinear form induced by the symmetric tensor $d_{ijk}$ on $\mathbb{C}^{8}$. Its finiteness follows from the finite number of nonvanishing structure constants of $\mathfrak{su}(3)$. While its precise numerical value is not needed for the formulation of Theorem 1, the definition (30) shows that $C_{3}^{(0)}$ (and hence also $C_{3}$ defined above) depends only on the algebraic data of $\mathfrak{su}(3)$ and is independent of the particular map $\Phi$. This ensures that the condition (15) is purely structural. In contrast, the quantity $c_{3}(\mu)$ from (14) depends on the Bloch parameters and captures how strongly the scalar part dominates in the Kadison–Schwarz expansion for the given diagonal map.